OpenStudy (anonymous):

Simplify each expression 5x - 3y - 2x + 3y 8x - 7y - (3 - 6y) 3(r - 10s) - 4(7s + 2r) 2(4z - 2x + y) - 4 (5z + x - y)

OpenStudy (anonymous):

5x - 3y - 2x + 3y =3x 8x - 7y - (3 - 6y) =8x-7y-3+6y =8x-y-3 3(r - 10s) - 4(7s + 2r) =3r-30s-28s-8r =-5r-58s 2(4z - 2x + y) - 4 (5z + x - y) =8z-4x+2y-20z-4x+47 =-8x+2y-12z+47

OpenStudy (anonymous):

1)3x 2.)8x-y-3 3)-5r-58s 4)-8+2y-12z+47